Solve using the five-step problem-solving process. Show all steps necessary to arrive at your solution.
The second angle of a triangle is 3 times as large as the first. The third angle is 55° more than the first. Find the measure of the smallest angle.

Answer by mananth(16949)

let first angle = x
second angle = 3x
third angle = x+55
..
The sum of angles of triangle = 180(deg)
x+3x+x+55=180
5x+55=180
add -55
5x+55-55=180-55
5x=125
divide by 5
5x/5=125/5
x=25
first angle 25 deg
second angle = 3x= 75 deg
third angle = x+55 = 25+55=80 deg.
CHECK
80+75+25=180